An airline records flight delays in and out of Chicago O'Hare airport for a year. The average delay for these flights is 9.85 minutes with a standard deviation of 3.27 minutes. For a sample of 58 flights, 16% of flights will have an average delay greater than how many minutes?

1) 13.1

2) There is not enough information to determine this.

3) 9.42

4) 6.6

5) 10.28

Please try and answer using a graphing calculator
Answer #1
mean μ= 9.85
standard deviation σ= 3.2700
sample size       =n= 58
std error=σ=σ/√n= 0.429

as top 16% flight times will fall at 84th percentile';

for 84th percentile critical value of z= 0.99
therefore corresponding value=mean+z*std deviation= 10.28 minutes